# Automated Contrastive Learning Strategy Search for Time Series

### v1

@@ 2.2. Solution Space
2.3. Search Algorithm
As shown in Section 2.2, there are around $3\times 10^{12}$ total strategies in the solution space. Therefore, it could be extremely challenging and resource consuming to manually pick a suitable CLS from the entire space. In this paper, we aim to automate the process of searching the suitable CLS within the solution space. In the following content, we first present an overview of the algorithmic part of AutoCL, and then elaborate its details, including the controller network and the two phases: candidate search and candidate evaluation.
Overview. In recent years, Reinforcement Learning (RL) has become a popular paradigm to implement AutoML pipelines. Our proposed algorithm follows the RL paradigms in (Cubuk et al., 2019; Li et al., 2022; Zhao et al., 2021), and an illustration is shown in Figure 2. The search algorithm is comprised of two phases: candidate search and candidate evaluation. In the candidate search phase, the agent, i.e., controller network $f_{C}$, interacts with the environment, including encoder training and validation, etc., for a certain number of iterations. For each interaction iteration, $f_{C}$ first samples a CLS $A$ and send it to the environment. Then, the environment trains the encoder $f_{E}$ based on $A$ and returns a reward $R$ to the agent. Finally, the agent updates the controller network $f_{C}$ based on $R$. During the interactions, a CLS $A$ satisfying certain requirements is recorded in the candidate set $\mathcal{A}$. In the candidate evaluation phase, for each CLS candidate $A\in\mathcal{A}$, we train an encoder $f_{E}$ based on $A$, then evaluate $f_{E}$ on the downstream task, and record its performance on the validation data.
Controller Network. The controller network $f_{C}$ learns to sample CLS for the given dataset and downstream task. An illustration of the controller network is presented in Figure 3. The controller network is comprised of (1) a learnable embedding $\mathbf{e}\in\mathbb{R}^{d}$, which learns the information of the input dataset and task during training; (2) a project function, projecting $\mathbf{e}$ into a hidden space; (3) $N$ action branches, each of which samples an option (e.g., 0.5) for its corresponding sub-dimension (e.g., point masking) in the solution space. Formally, $a_{n}$ is sampled according to the probability $\mathbf{p}_{n}\in\mathbb{R}^{n_{K}}$:
(1) $\mathbf{p}_{n}=\text{MLP}_{n}(\sigma(\mathbf{W}\mathbf{e})),$
where $\mathbf{W}\in\mathbb{R}^{d\times d}$ and $\sigma$ are the weight and the tanh activation of the project layer. We denote a full strategy as $A=[a_{1},a_{2},\dots,a_{N}]$.
Figure 3. Controller network $f_{C}$.
Candidate Search. During candidate search, parameters of both the controller $f_{C}$ and the encoder $f_{E}$ need to be updated. We denote the parameters of $f_{C}$ and $f_{E}$ as $\theta_{C}$ and $\theta_{E}$, respectively. Optimizing $\theta_{C}$ and $\theta_{E}$ is a bi-level optimization problem (Zhao et al., 2021; Li et al., 2022; Luo et al., 2023). Given the strategy sampled by the controller $A\sim f_{C}$, the lower-level problem is to find the optimal $\theta_{E}^{*}$ that minimizes the contrastive loss $L\in A$. The upper-level problem is to maximize the reward $R$ on the validation set by optimizing $\theta_{C}$. Formally, we have:
@@ 2.3. Search Algorithm
In practice, optimizing $\theta_{C}$ after the completion of optimizing $\theta_{E}$, i.e., conducting the complete contrastive pre-training for $f_{E}$, could be extremely costly (Zhao et al., 2021; Li et al., 2022). Following (Liu et al., 2018), we use the first-order approximation of the gradient for the encoder $f_{E}$. Let $\eta$ be the learning rate in gradient decent, then $\theta_{E}^{*}(A)$ is approximated by:
@@
In the bi-level optimization, if the model in the lower-level optimization is trained along an undesired direction, it will take much more steps to bring it back to the desired direction (Li et al., 2022). To address this issue, in the lower-level optimization, rather than directly updating the parameters of the encoder $f_{E}$, we update the parameters of $\tilde{f}_{E}$, which is the copy of $f_{E}$. Let $\tilde{f}_{E}^{*}$ denote the updated encoder copy with parameters $\tilde{\theta}_{E}^{*}$. Then we can obtain the raw reward $R(A,\tilde{\theta}_{E}^{*})$ and the final reward $\Delta$ based on $\tilde{f}_{E}^{*}$. If $\Delta<0$, then we discard $\tilde{f}_{E}^{*}$ since it goes to the undesired direction. If $\Delta>0$, then we replace the original encoder $f_{E}$ with $\tilde{f}_{E}^{*}$, and also add $A$ to the candidate set $\mathcal{A}$, as the model is updated in the desired direction.
@@
Candidate Evaluation. To evaluate each candidate $A\in\mathcal{A}$, we first pre-train the encoder $f_{E}$ on the training data based on $A$. Then we train a full model $f=f_{D}\circ f_{E}$ on the training data for the downstream task (e.g., classification), where $f_{D}$ is the downstream model. The evaluation score (e.g., accuracy) of $f$ on the validation data is recorded as the performance of $A$.
@@
Complexity Analysis. Our proposed algorithm is efficient in both phases. In the phase 1, we only use the first order approximation rather than exhausted training; in the phase 2, we could take advantage of parallel computing. Denote $N_{cs}$ as the number of iterations in phase 1, $T_{ep}$ and $T_{val}$ as the time of training the encoder for one epoch, and the time of validation. The time complexity for the phase 1 is $O(N(T_{ep}+T_{val}))$. Denote $M$ as the number of iterations of the full contrastive pre-training, $|\mathcal{A}|$ as the total number of candidates, and $K$ as the number of machines available for evaluation, then the time complexity of the phase 2 is $O(\frac{|\mathcal{A}|}{K}(MT_{ep}+T_{val}))$. Therefore, the total complexity is $O((\frac{|\mathcal{A}|}{K}M+N_{cs})T_{ep}+(N_{cs}+1)T_{val})$ In comparison, if we do not use the first order approximation in Equation (3), then the complexity is $O(N_{cs}MT_{ep}+N_{cs}T_{val})))$ for phase 1, which is significantly larger than our current practice. Note that $N_{cs}\geq|\mathcal{A}|$.
@@
3. Experiments
In this section, we provide empirical evaluations to answer the questions: Q1. How well does the CLS found by AutoCL perform compared to existing methods? Q2. Does there exist a Generally Good Strategy (GGS) that has a strong performance on all settings? Q3. How does each component in AutoCL contribute to the final results? Q4. What can we learn from the candidates found by AutoCL? Q5. How does AutoCL perform in the real-world applications? We answer these questions in Section 3.2 - 3.6, respectively.
3.1. Experimental Setup
Datasets. We use 7 real-world public datasets and 1 private data. The public offline datasets can be grouped by three tasks: classification, anomaly detection and forecasting. For classification: we use HAR (Anguita et al., 2013), Epilepsy (Andrzejak et al., 2001; Eldele et al., 2021). HAR is comprised of 10,299 sensor readings collected from 30 subjects, where each reading corresponds to one of the 6 activities and has a length of 128. The number of samples in the training, validation and test data are 5,881/1,471/2,947. Epilepsy contains 11,500 electroencephalogram (EEG) time series segments in a length of 178 from 500 subjects, and the task is to recognize whether there is epileptic seizure. The train/validation/test split is 7,360/1,840/2,300. For anomaly detection, we use Yahoo (Laptev et al., 2015) and KPI (Ren et al., 2019). Yahoo has 367 hourly time series with human labeled anomaly points, which contains synthetic and real web traffic data to Yahoo! services. KPI is comprised of 58 minutely KPI curves collected from various Internet companies such as Tencent and eBay. The two anomaly detection datasets are originally split into 50%/50% for train/test. During searching, we further split the training data by 90%/10% for training and validation. For forecasting, we use ETTh1/h2/m1(Zhou et al., 2021), which are 2-year time series data of oil temperature of electricity transformers. ETTh1 and ETTh2 are hourly recordings, which contain 17,420 time steps in total, and ETTm1 is a 15 min recording, which contains 69,680 time steps. These datasets are split into 12/4/4 months for train/validation/test. The private online dataset contains SEEG (stereo-electroencephalogram) signals of 15 anonymized patients collected from a the epilepsy surgery department of a Chinese hospital - Ruijin Hospital. For each patient, 12-15 invasive electrodes each with 12-18 contacts is used to obtain about 3.5 hour SEEG readings. Two neurosurgery technicians jointly marked the onset and end times of the seizures. We follow (Chen et al., 2022; Andrzejak et al., 2001) to pre-process the dataset. The signal frequency is unified to 256Hz, and the time series are divided into segments length of 0.5s. The segments containing epileptic waves are considered as positive samples accounting for 1.6% of the data. Following (Andrzejak et al., 2001), we balance the two classes into 50%:50%, which results in 11,778 samples. We further split the data into 70%/10%/20% for train/validation/test for each patient.
Evaluations and Implementations. Please refer to Appendix C for evaluation metrics, baseline settings and implementation details.
3.2. Direct Application of AutoCL
In this subsection, we answer Q1 by demonstrating the effectiveness of AutoCL for searching the suitable CLS for a given dataset and task. The overall performance of different methods on 7 public benchmark datasets and 3 downstream tasks are presented in Table 2-4. Among all the baselines, InfoTS, which could automatically search the suitable data augmentations, performs better than other baselines, which use handcrafted CLS. This observation shows the effectiveness of automatically searching the data augmentations. AutoCL further outperforms InfoTS, indicating the overall effectiveness of the proposed search space and search algorithm.
@@ 3.3. Transferability Study
Although AutoCL could efficaciously discover the suitable CLS for given datasets/tasks, in many cases, searching could still be resource consuming. Therefore, we seek to find a Generally Good Strategy (GGS) (Q2) which can be used as a strong baseline and a good starting point for new datasets and tasks.
We search GGS on 3 datasets HAR/Yahoo/ETTh1, and then test it on all the datasets. The procedure of finding GGS is described as follows. (1) Collect sets of top $K$ candidate CLS from each of HAR/Yahoo/ETTh1, and use Cartesian product to obtain the $K^{3}$ pairs $(A_{1},A_{2},A_{3})$. (2) Obtain the top candidate pair by the number of the shared sub-dimensions across $A_{1},A_{2},A_{3}$, and add the shared options in GGS. (3) For the un-shared sub-dimensions, discard the options that lead to significant validation performance drop. For details, please refer to appendix.
@@
A GGS we found is shown in Table 5. There are some interesting findings for the 4 dimensions. For the data augmentations, small perturbations on the input time series are preferred. For the embedding transformations, large jitterring is preferred. Besides, as a compromise for generally good across different datasets and tasks, normalization is not adopted by GGS. In fact, we have observed in Section 3.5 that classification tasks could benefit from LayerNorm yet forecasting disfavors normalization. For the contrastive pair construction, the instance contrast is the most important while the other aspects, i.e., the temporal and cross-scale contrasts, are less important. For the contrastive losses, the popular InfoNCE loss is generally good for all datasets and tasks. For the similarity function, negative Euclidean distance is preferred over the popular dot product and cosine similarity. Euclidean distance measuress the distance between two point, yet cosine similarity and dot product also consider the angle between the two points. This observation indicates that distance is more effective than angles for embeddings of time series data.
The performance of GGS on all datasts/tasks is presented in Table 2-4. It can be observed that GGS outperforms all the baselines in general, but underperforms AutoCL. For practical usage, we leave the trade-off, between AutoCL from scratch (with better performance) and direct transferring GGS (with lower computational cost) to novel settings, to the readers.
@@
3.4. Ablation Study
To answer Q3, we investigate the impact of different components of AutoCL on HAR, and the results are shown in Table 6.
Effectiveness of the solution space. Compared with existing studies, which only consider the data augmentation dimension of CLS, our proposed solution space is more comprehensive, which is comprised of 4 most important dimensions: data augmentations, embedding transformations, contrastive pair construction and contrastive losses (Table 1). We can observe that the CLS obtained from the full space significantly outperforms the CLS obtained from the data augmentation dimension alone.
Effectiveness of the first-order approximation and the separate search/evaluation phases. We compare AutoCL with the full pre-training of $f_{E}$ during phase 1, i.e., given the CLS $A$ sampled by $f_{C}$, we pre-train the encoder $f_{E}$ from scratch till the max number of iterations is reached. The results in Table 6 show that the CLS found by the first-order approximation (AutoCL) is competitive to the full pre-training of $f_{E}$ in terms of ACC and F1. However, AutoCL is more than 8$\times$ faster than the full pre-training. This is because AutoCL (1) only pre-trains $f_{E}$ for one epoch for each iteration in phase 1; (2) evaluates candidates in parallel in phase 2.
3.5. Empirical Analysis of the Candidates
To answer Q4, we provide in-depth empirical analysis based on the candidate CLS, and hope to provide potential insights and guidelines for future CLS designs. Violin plots of 6 sub-dimensions for HAR/Yahoo/ETTh1 are shown in Figure 5, where rows correspond to datasets and columns correspond to sub-dimensions.
Normalization. The 4th column shows that LayerNorm is important for classification, however, the embedding normalization has negative impact for forecasting and anomaly detection. We suspect this is because the normalization might destroy the fine-grained information, e.g., small anomaly points, which might be critical for forecasting and anomaly detection.
Embedding Jittering. The 5th column indicates that applying jittering on embeddings is generally useful. This finding is consistent with (Wang et al., 2019) that jittering injects various small semantic perturbations into the embeddings, which helps the model to capture the core semantics during contrastive pre-training.
Frequency Masking. In the last column, Figure 4(r) shows that frequency masking with high masking ratio could easily lead to good performance (low MSE). We believe this is because frequency masking could remove the short-range noisy patterns, and make the model focus on the long-range patterns. However, Figure 4(l) shows that the frequency masking has negative impact on anomaly detection. Essentially, the anomaly points are sudden and short-range noise, which only accounts for a small ratio, e.g., 1% of the entire time series. Fourier Transform (FT) and Inverse Fourier Transform (IFT) involved in the frequency masking can easily remove these points from the input, and thus the model is unable to learn the information about anomaly points during pre-training.
Table 7. Classification results in deployed application.
Metrics AutoCL GGS+LayerNorm GGS InfoTS TS2Vec
ACC 0.760 0.739 0.737 0.735 0.728
F1 0.756 0.734 0.730 0.730 0.719
3.6. Experiments in Deployed Application
To answer Q5, we deploy AutoCL to the machine learning platform of our partner hospital. The medical practitioners can easily utilize our AutoCL to obtain the effective time series models for analyzing various physiological time series data collected in medicine, e.g., EEG and SEEG time series signals in epilepsy healthcare.
@@ 3.6. Experiments in Deployed Application
Figure 5. Violin plots of 6 sub-dimensions for HAR/Yahoo/ETTh1. X-axis: options. Y-axis: ACC/F1/MSE(48 horizon). ACC/F1: the higher the better. MSE: the lower the better. Low, Medium and High correspond to $0.1\sim 0.3$, $0.4\sim 0.6$ and $0.7\sim 0.95$ for embedding jittering and frequency masking ratio.
4. Related Work
Time Series Contrastive Learning. Contrastive learning aims to train encoders that can learn informative embeddings by distinguishing positive data pairs and negative data pairs (Liu et al., 2021), and it has been demonstrated to be effective in domains such as computer vision (Ozbulak et al., 2023), natural language processing (Zhang et al., 2022a) and graph mining (Liu et al., 2022). In recent years, some studies introduce contrasive learning into the time series domain and have achieved impressive performance (Zhang et al., 2023). SRL (Franceschi et al., 2019) proposes to treat two adjacent sub-sequences from the same time series as positive pairs, and sub-sequences from other time series as negative samples. CPC (Oord et al., 2018) introduces the popular InfoNCE loss. Self-EEG (Sarkar and Etemad, 2020) augments the input EEG signals with various transformations. TNC (Tonekaboni et al., 2020) constructs contrastive pairs based on the local smoothness of time series. TS-TCC (Eldele et al., 2021) transforms the input data with weak and strong augmentations, and use temporal and contextual contrasting to train the encoders. Instead of only considering data augmentations in the temporal domain, CoST (Woo et al., 2021), BTSF (Yang and Hong, 2022) and TF-C (Zhang et al., 2022b) further propose to augment time series in the frequency domain. TS2Vec (Yue et al., 2022) obtains time series representations of multiple scales by hierarchical pooling and performs CL at each scale. To reduce human efforts in finding the optimal data augmentations, InfoTS (Luo et al., 2023) recently proposes to automatically select data augmentations through meta-learning. However, InfoTS only considers data augmentations yet ignores other critical components in contrastive learning such as embedding transformation, contrasive pair construction and loss functions. Besides, InfoTS selects data augmentations based on a dataset/task-agnostic criteria, which is less effective in capturing the characteristics of the datasets and downstream tasks and thus might not be able find the optimal solution. Compared with InfoTS, AutoCL  (1) considers a larger search space comprised of data augmentations, emebedding transformations, and loss functions; (2) finds the optimal strategy based on the model’s performance on the validation set.
Automated Contrastive Learning. Despite various powerful contrastive strategies have been developed for particular datasets and downstream tasks, they cannot be easily adapted to other datasets and tasks (Luo et al., 2023; Tian et al., 2020; You et al., 2021). Therefore, it is necessary to develop automatic searching algorithms to find the optimal contrastive strategies for a given dataset and task. In the computer vision domain, InfoMin (Tian et al., 2020) introduces a mutual information minimization criteria for selecting augmented views. SelfAugment (Reed et al., 2021) uses the self-supervised image rotation task to guide the selection of data augmentations. In the graph domain, JOAO (You et al., 2021) presents a Bayesian framework to adaptively find the optimal data augmentations. AutoSSL (Jin et al., 2021) learns the weights of different contrastive strategies. AutoGCL (Yin et al., 2022) builds an end-to-end framework to automatically learn the distributions of node level augmentations. In the time series domain, InfoTS (Luo et al., 2023) recently propose to adaptively select the optimal data augmentations based on information theory. Different from previous studies which mainly focus on searching the optimal data augmentations, in this paper, we propose AutoCL to simultaneously consider various dimensions of CL, including data augmentations, embedding transformations, contrastive pair construction and loss functions.

### v2

@@ 2.3. Search Algorithm
Overview. In recent years, Reinforcement Learning (RL) has become a popular paradigm to implement AutoML pipelines. Our proposed algorithm follows the RL paradigms in (Cubuk et al., 2019; Li et al., 2022a; Zhao et al., 2021), and an illustration is shown in Figure 2. The search algorithm has two phases: candidate search and candidate evaluation. During candidate search, the agent, i.e., controller network $f_{C}$, interacts with the environment, including encoder training and validation, etc., for several iterations. For each iteration, $f_{C}$ samples a CLS $A$ and sends it to the environment. Then, the environment trains the encoder $f_{E}$ based on $A$ and returns a reward $R$ to $f_{C}$. Finally, $f_{C}$ is updated based on $R$, and $A$ satisfying certain requirements is recorded in the candidate set $\mathcal{A}$. During candidate evaluation, for each candidate $A\in\mathcal{A}$, we train $f_{E}$ based on $A$, then evaluate $f_{E}$ on the downstream task and record its performance on the validation data.
Controller Network. The controller network $f_{C}$ learns to sample CLS for the given dataset and downstream task, an illustration of which is shown in Figure 3. $f_{C}$ is comprised of (1) a learnable embedding $\mathbf{e}\in\mathbb{R}^{d}$, which learns the information of the given dataset and task; (2) a project function, projecting $\mathbf{e}$ into a hidden space; (3) $N$ action branches, each of which samples an option, e.g., 0.5, for its corresponding sub-dimension, e.g., point masking. Formally, $a_{n}$ is sampled according to the probability $\mathbf{p}_{n}=\text{MLP}_{n}(\sigma(\mathbf{W}\mathbf{e}))\in\mathbb{R}^{n_{K}}$, where $\mathbf{W}\in\mathbb{R}^{d\times d}$ and $\sigma$ are the weight and the tanh activation of the project layer. We denote a full strategy as $A=\{a_{1},a_{2},\dots,a_{N}\}$.
Candidate Search. During candidate search, parameters of both the controller $f_{C}$ and the encoder $f_{E}$ need to be updated. Optimizing their parameters $\theta_{C}$ and $\theta_{E}$ is a bi-level optimization problem (Zhao et al., 2021):
@@
Given the strategy sampled by the controller $A\sim f_{C}$, the lower-level problem is to find the optimal $\theta_{E}^{*}$ that minimizes the contrastive loss $L\in A$. The upper-level problem is to maximize the reward $R$ on the validation set by optimizing $\theta_{C}$. In practice, optimizing $\theta_{C}$ after the completion of optimizing $\theta_{E}$, i.e., conducting the complete CL for $f_{E}$, can be extremely costly (Zhao et al., 2021; Li et al., 2022a). Following (Liu et al., 2018), we use the first-order approximation of the gradient for $f_{E}$. Let $\eta$ be the learning rate in gradient decent, then $\theta_{E}^{*}(A)$ is approximated by:
@@
In the bi-level optimization, if $\theta_{E}$ in the lower-level optimization is trained along an undesired direction, it will take many more steps to bring it back to the desired direction (Li et al., 2022a). To address this issue, rather than directly updating $\theta_{E}$, we make a copy of $\theta_{E}$, denoted by $\tilde{\theta}_{E}$, and update the copied parameters to $\tilde{\theta}_{E}^{*}$. Then we can obtain the raw $R(A,\tilde{\theta}_{E}^{*})$ and final $\Delta$ rewards based on $\tilde{\theta}_{E}^{*}$. If $\Delta<0$, then we discard $\tilde{\theta}_{E}^{*}$ since it goes in the undesired direction. If $\Delta>0$, then we replace the original encoder $\theta_{E}$ with $\tilde{\theta}_{E}^{*}$, and also add $A$ to the candidate set $\mathcal{A}$, as $\theta_{E}$ is updated in the desired direction.
@@
Candidate Evaluation. To evaluate each candidate $A\in\mathcal{A}$, we first pre-train the encoder $f_{E}$ on the pre-training data based on $A$. Then we train a full model $f=f_{D}\circ f_{E}$ on the training data of the downstream task, e.g., classification, where $f_{D}$ is the downstream model. The evaluation score, e.g., accuracy, of $f$ on the validation data is recorded as the performance of $A$.
@@
Complexity Analysis. AutoCL is efficient in both phases. In phase 1, we use the 1st-order approximation rather than exhausted training; in phase 2, we can take advantage of parallel computing. Let $N$ be the number of iterations in phase 1, $T_{t}$, be the time of training the encoder for one epoch, and $T_{v}$ be the time of validation. The time complexity for phase 1 is $O(NT_{t}+(N+1)T_{v})$. Let $M$ be the number of iterations of the full contrastive pre-training, $|\mathcal{A}|$ be the total number of candidates, and $K$ be the number of machines, then the time complexity of phase 2 is $O(\frac{|\mathcal{A}|}{K}(MT_{t}+T_{v}))$. Therefore, the total complexity is $O((N+\frac{|\mathcal{A}|}{K}M)T_{t}+(N+1+\frac{|\mathcal{A}|}{K})T_{v})$. In comparison, if we do not use the 1st order approximation in Equation (2), then the complexity is $O(NMT_{t}+(N+1)T_{v})$ for phase 1, which is significantly larger than our current practice.
3. Experiments
3.1. Experimental Setup
Datasets. We use 7 public and 1 private datasets. The public datasets can be grouped by three tasks: classification, anomaly detection and forecasting. For classification: we use HAR (Anguita et al., 2013), Epilepsy (Andrzejak et al., 2001; Eldele et al., 2021). HAR contains 10,299 sensor readings with a length of 128, where each reading corresponds to one of the 6 activities. Epilepsy contains 11,500 electroencephalogram (EEG) time series segments with a length of 178 from 500 subjects, and the task is to recognize epileptic seizures. For anomaly detection, we use Yahoo (Laptev et al., 2015) and KPI (Ren et al., 2019). Yahoo has 367 hourly time series with human-labeled anomaly points, which is web traffic data to Yahoo! services. KPI contains 58 minutely KPI curves from various Internet companies. These two datasets are originally split into 50%/50% for train/test. During searching, we further split the training data by 90%/10% for training and validation. For forecasting, we use ETTh1/h2/m1(Zhou et al., 2021), which are 2-year time series of oil temperature of electricity transformers. The private dataset contains SEEG (stereo-electroencephalogram) signals of 15 anonymized patients collected from Ruijin Hospital. For each patient, 12-15 invasive electrodes each with 12-18 contacts are used to obtain about 3.5 hour readings. Two neurosurgery technicians jointly marked the onset and end times of the seizures. The signal frequency is unified to 256Hz, and the time series are divided into segments length of 0.5s. The segments containing epileptic waves are considered as positive samples accounting for 1.6% of the data. Following (Andrzejak et al., 2001), we balance the two classes into 50%:50%, resulting in 11,778 samples. We split the data into 70%/10%/20% for train/validation/test for each patient.
Implementation. Most settings follow (Yue et al., 2022). The embedding size of $f_{C}$ is 320, and MLPs are linear layers with softmax activations.
3.2. Direct Application of AutoCL
In this subsection, we compare the proposed AutoCL with existing CL methods to demonstrate the effectiveness of AutoCL. The overall performance of different methods on 7 public benchmark datasets and 3 downstream tasks are presented in Table 2-4. Among all the baselines, InfoTS, which could automatically search suitable data augmentations, performs better than other baselines, which use handcrafted CLS. This observation shows the effectiveness of automatically searching the data augmentations. AutoCL further outperforms InfoTS, indicating the overall effectiveness of the proposed search space and search algorithm.
@@ 3.3. Transferability Study
Although AutoCL could efficaciously discover the suitable CLS for given datasets/tasks, in many cases, searching could still be resource-consuming. Therefore, we seek to find a Generally Good Strategy (GGS) which can be used as a strong baseline and a good starting point for new datasets and tasks.
We search GGS on 3 datasets HAR/Yahoo/ETTh1, and then test it on all the datasets. The procedure of finding GGS is described as follows. (1) Collect sets of top $K$ candidate CLS from each of HAR/Yahoo/ETTh1, and use Cartesian product to obtain the $K^{3}$ pairs $(A_{1},A_{2},A_{3})$. (2) Obtain the top candidate pair by the number of the shared sub-dimensions across $A_{1},A_{2},A_{3}$, and add the shared options in GGS. (3) For the un-shared sub-dimensions, discard the options that lead to a significant validation performance drop.
@@
A GGS we found is shown in Table 5. There are some interesting findings for the 4 dimensions. For the data augmentations, small perturbations on the input time series are preferred. For the embedding transformations, large jitterring is preferred. Besides, as a compromise for the general good across different datasets and tasks, normalization is not adopted by GGS. In fact, we have observed in Section 3.5 that classification tasks could benefit from LayerNorm yet forecasting disfavors normalization. For the contrastive pair construction, the instance contrast is the most important while the other aspects, i.e., the temporal and cross-scale contrasts, are less important. For the contrastive losses, the popular InfoNCE loss is generally good for all datasets and tasks. For the similarity function, negative Euclidean distance is preferred over the popular dot product and cosine similarity. Euclidean distance measures the distance between two points, yet cosine similarity and dot product also consider the angle between the two points. This observation indicates that distance is more effective than angles for embeddings of time series data. The performance of GGS on all datasets/tasks is presented in Table 2-4. It can be observed that GGS outperforms all the baselines in general, but underperforms AutoCL.
3.4. Ablation Study
In this subsection, we investigate the impact of different components of AutoCL on HAR, and the results are shown in Table 6.
Effectiveness of the solution space. Compared with existing studies, which only consider the data augmentation dimension of CLS, our proposed solution space is more comprehensive, which is comprised of 4 most important dimensions: data augmentations, embedding transformations, contrastive pair construction and contrastive losses (Table 1). We can observe that the CLS obtained from the full space (AutoCL) significantly outperforms the CLS obtained from the data augmentation dimension alone (Data Aug. Only).
Effectiveness of the first-order approximation and the separate search/evaluation phases. We compare AutoCL with the full pre-training of $f_{E}$ during phase 1, i.e., given the CLS $A$ sampled by $f_{C}$, we pre-train the encoder $f_{E}$ from scratch till the max number of iterations is reached. The results in Table 6 show that the CLS found by the first-order approximation (AutoCL) is competitive to the full pre-training of $f_{E}$ in terms of ACC and F1. However, AutoCL is more than 8$\times$ faster than the full pre-training. This is because AutoCL: (1) only pre-trains $f_{E}$ for one epoch for each iteration in phase 1; (2) evaluates candidates in parallel in phase 2.
@@ 3.4. Ablation Study
Figure 5. Violin plots of 6 sub-dimensions for HAR/Yahoo/ETTh1. X-axis: options. Y-axis: ACC/F1/MSE(48 horizon). ACC/F1: the higher the better. MSE: the lower the better. Low, Medium and High correspond to $0.1\sim 0.3$, $0.4\sim 0.6$ and $0.7\sim 0.95$.
3.5. Empirical Analysis of the Candidates
In this subsection, we provide in-depth empirical analysis based on the candidate CLS, and hope to provide potential insights and guidelines for future CLS designs. Violin plots of 6 sub-dimensions for HAR/Yahoo/ETTh1 are shown in Figure 5, where rows correspond to datasets and columns correspond to sub-dimensions.
Normalization. The 4th column shows that LayerNorm is important for classification, however, the embedding normalization has a negative impact on forecasting and anomaly detection. We suspect this is because the normalization might destroy the fine-grained information, e.g., small anomaly points, which might be critical for forecasting and anomaly detection.
Embedding Jittering. The 5th column indicates that applying jittering on embeddings is generally useful. This finding is consistent with (Wang et al., 2019) that jittering injects various small semantic perturbations into the embeddings, which helps the model capture the core semantics during contrastive pre-training.
Frequency Masking. In the last column, Figure 5(r) shows that frequency masking with a high masking ratio could easily lead to good performance (low MSE). We believe this is because frequency masking could remove the short-range noisy patterns, and make the model focus on the long-range patterns. However, Figure 5(l) shows that frequency masking has a negative impact on anomaly detection. Essentially, the anomaly points are sudden and short-range noise, which only accounts for a small ratio, e.g., 1% of the entire time series. Fourier Transform (FT) and Inverse Fourier Transform (IFT) involved in the frequency masking can easily remove these points from the input, and thus the model is unable to learn the information about anomaly points during pre-training.
3.6. Experiments in Deployed Application
We deploy AutoCL to the machine learning platform of our partner, Shanghai Ruijin Hospital, a top-tier hospital in China. The medical practitioners can easily utilize our AutoCL to obtain effective time series models for analyzing various physiological time series data collected in medicine, e.g., EEG and ECG time series in healthcare.
Table 7. Classification results on the real-world application.
Metrics AutoCL GGS+LayerNorm GGS InfoTS TS2Vec
ACC 0.760 0.739 0.737 0.735 0.728
F1 0.756 0.734 0.730 0.730 0.719